OpenStudy (anonymous):

integral of 1/x(sqrtx)

OpenStudy (anonymous):

(-2)*1/sqrt(x)

OpenStudy (anonymous):

how did you get that?

OpenStudy (shamim):

the formula u hv to use integral x^n=(x^n+1/n+1)

OpenStudy (anonymous):

so a basic u sub?

OpenStudy (anonymous):

yes, and you have x^(-3/2), so the answer is x^(-3/2+1)/(-3/2+1)=x^(-1/2)/(-1/2)=(-2)/sqrt(x)

OpenStudy (ipwnbunnies):

Yeah, the trick is to turn that square root into an exponent, so you can apply the Power Rule.

OpenStudy (anonymous):

thanks

OpenStudy (shamim):

welcome